here's what my teacher sent us as an assignment for our topic which is function,
if f(x) = (x^2) - (3x^3), then find x+h.
i just don't know what to do first. what is the answer and how will i answer? thanks!

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

if f(x) = (x^2) - (3x^3), then find x+h.
------------
Wherever there's an x, sub x+h
That's all.
f(x+h) = (x+h)^2 - 3(x+h)^3
-------
You can expand and simplify.

Answer by solver91311(24713)



Write out in pencil. Then erase every and replace each one with . Then simplify the resulting expression.

The results of your first step should look like:



Do the binomial expansions and then collect like terms.

John

My calculator said it, I believe it, that settles it